You just purchased a share of SPCC for $96. You expect to receive a dividend of ​$3 in one year. If you expect the price after the dividend is paid to be $116​, what total return will you have earned over the​ year? What was your dividend​ yield? Your capital gain​ rate?
Answer #1

Total return =(116-96+3)/96 =23.96%

Dividend yield =3/96 =3.13%

Capital gain yield =Total return-dividend yield =20.83%